### v1

Abstract
The integration of Large Language Models (LLMs) like GPT-4 into traditional Natural Language Processing (NLP) tasks has opened new avenues for enhancing model performance while reducing the reliance on extensive human annotations. This paper presents a novel approach that leverages the Chain of Thought (CoT) prompting technique to distill knowledge from GPT-4, subsequently applying it to improve the efficiency and effectiveness of a smaller model, BERT, on Named Entity Recognition (NER) tasks. Our method involves a two-phase training process: initially employing GPT-4 annotated data for pre-training and then refining the model with a combination of distilled and original human-annotated data. The results demonstrate that our mixed-training strategy significantly outperforms models trained solely on human annotations, achieving superior F1-scores and showcasing a cost-effective solution for resource-limited or closed-network settings. The study also discusses the challenges encountered, such as LLM output variability and the tendency towards hallucinations, proposing future work directions to enhance prompt design and annotation selection. Our findings indicate a promising synergy between LLM insights and traditional NLP techniques, paving the way for more accessible and robust NLP applications.
Introduction
The advent of Large Language Models (LLMs) such as the Generative Pre-trained Transformer (GPT) series has marked a significant milestone in the field of artificial intelligence and natural language processing (NLP). With their remarkable reasoning capabilities and extensive knowledge bases, LLMs have demonstrated unparalleled performance across a wide range of applications, from text generation and dialogue systems to content comprehension. Despite their success, these models face a notable performance gap in traditional NLP tasks like Named Entity Recognition (NER) and Relation Extraction (RE) when compared to the current state-of-the-art (SOTA) technologies. However, with continuous advancements, there is a promising potential for LLMs to not only bridge this gap but also to significantly reduce costs and enhance efficiency in traditional NLP tasks.
The rising dominance of LLMs has sparked debates regarding the relevance of traditional NLP techniques. Some argue that the capabilities of LLMs might render traditional methods obsolete. Yet, this perspective overlooks the intrinsic limitations of LLMs, such as their capacity constraints and tendency towards hallucination—producing outputs that may diverge from factual accuracy. In reality, traditional NLP techniques remain indispensable, especially in applications such as knowledge graph construction where tasks like NER and RE are crucial for extracting structured information from unstructured text. Therefore, LLMs and traditional NLP techniques are not in competition but rather complement each other, forming a synergistic relationship that leverages the strengths of both approaches.
However, integrating LLMs into traditional NLP tasks is not without challenges. First, the inherent differences in task nature between generative models like LLMs and the sequence labeling or classification tasks of traditional NLP present a significant gap. Bridging this gap requires innovative approaches that have yet to be fully explored. Additionally, LLMs are prone to producing "hallucinations," which can compromise the accuracy of generated annotations. This issue is exacerbated by the dependency of output quality on the design of prompts, introducing variability and potential inaccuracies in task execution. Moreover, the substantial computational resources required for LLMs pose a barrier to their application in resource-constrained environments or scenarios with stringent privacy and security requirements, necessitating solutions that can operate within these limitations.
Addressing these challenges, our research proposes a comprehensive solution utilizing knowledge distillation techniques. By leveraging LLMs for data annotation and incorporating these annotations in the training of smaller models like BERT, we enhance the performance of these models on traditional NLP tasks. Our approach mitigates the limitations of LLMs through the use of Chain of Thought (CoT) prompting to refine reasoning processes and generate more accurate annotations. Furthermore, we introduce an efficient prompt design that effectively translates LLM outputs into usable formats for sequence labeling tasks, overcoming issues of entity overlap and discontinuation common in many domains. Through a series of experiments utilizing datasets such as CONLL and Xtreme, we demonstrate the effectiveness of our mixed training strategy, which combines distilled LLM annotations with manually annotated data, achieving superior performance over models trained solely on human-annotated datasets.
This framework not only validates the feasibility of enhancing traditional NLP tasks with LLM-generated annotations but also showcases a scalable methodology that can be adapted across various LLMs, smaller models, and datasets. Our findings underscore the potential of this approach to reduce manual annotation costs, increase efficiency, and enable the deployment of capable models in resource-limited and privacy-sensitive settings, thereby filling a critical gap left by the limitations of LLMs in these contexts.
Related works
The integration of Large Language Models (LLMs) into Named Entity Recognition (NER) tasks has seen varied approaches and methodologies, demonstrating the evolving landscape of NLP research in leveraging advanced AI models for enhancing traditional tasks. This section outlines related work that has contributed to the development of methods and strategies in applying LLMs and other advanced techniques to NER, highlighting the diverse efforts to address challenges such as data scarcity, model efficiency, and cross-lingual applications.
Cross-lingual NER remains a significant challenge due to the scarcity of annotated training data, especially in low-resource languages. To address this, [1] introduces a Cross-lingual Entity Projection framework (CROP) that leverages a multilingual labeled sequence translation model for zero-shot cross-lingual NER, demonstrating substantial improvements over previous approaches. The gap between the sequence labeling nature of NER tasks and the text-generation orientation of LLMs is explored in [2], where the GPT-NER model is proposed. By transforming the NER task into a text generation format, GPT-NER addresses the hallucination issue common in LLMs, showing promise in low-resource and few-shot settings. PromptNER [3] presents a novel few-shot and cross-domain NER algorithm that utilizes prompt-based heuristics with LLMs, requiring only a few-shot examples and entity type definitions to achieve state-of-the-art performance on several benchmarks. This underscores the potential of prompt-based methods in NER tasks. [4] explores targeted distillation with mission-focused instruction tuning to train student models for NER tasks, demonstrating that these models can significantly outperform general instruction-tuned models across a broad application class, including open information extraction. The Multi-view Contrastive Learning approach for Cross-lingual NER (MCL-NER) proposed in [5] tackles the challenge of semantic and token-level representation alignment across languages, achieving new state-of-the-art performance on the XTREME benchmark for 40 languages. Inspired by the Chain of Thought methodology, [6] leverages the step-by-step problem-solving capabilities of LLMs for Biomedical NER, incorporating external knowledge to enhance entity type determination, showcasing significant improvements in a domain-specific application. [8] focuses on zero-shot information extraction in NER, adapting reasoning strategies tailored for NER tasks with LLMs. The introduction of syntactic augmentation and a two-stage majority voting strategy demonstrates remarkable improvements across multiple benchmarks. Lastly, [9] addresses the issue of applying pre-trained transformer models to long documents for NER tasks. By generating a synthetic context retrieval training dataset and training a neural context retriever based on a BERT model, this work shows that it is possible to outperform unsupervised retrieval baselines, marking an advancement in handling long-document NER tasks.
These works collectively underscore the dynamic interplay between LLMs, traditional NLP techniques, and innovative methodologies in advancing the field of NER. Our research builds upon these foundations, aiming to further leverage the capabilities of LLMs to enhance NER model training and performance, particularly through the integration of Chain of Thought prompting and knowledge distillation techniques.
Methodology
Overview
This study investigates the potential of enhancing traditional Natural Language Processing (NLP) tasks through the integration of Large Language Models (LLMs), specifically using GPT-4 for annotation and BERT for model training. Focusing on the Named Entity Recognition (NER) task within the widely recognized CONLL2003 benchmark, our approach is divided into two main phases. In the first phase, GPT-4 is employed to automatically annotate a selected subset of the dataset without any fine-tuning, aiming to assess the feasibility and quality of LLM-generated annotations. The second phase examines the impact of these LLM annotations on the performance of the BERT model, which is trained using a mix of original and LLM-annotated data. This comparative analysis seeks to uncover the benefits and limitations of leveraging LLM annotations in the context of traditional NLP frameworks, with the ultimate goal of improving efficiency and accuracy in language processing tasks.
Phase One: Data Annotation with LLM
Benchmark Dataset
Our study employs the CONLL2003 dataset, a cornerstone in the field of named entity recognition, focusing on four types of named entities: persons, locations, organizations, and miscellaneous entities that do not fit into the aforementioned categories.
For the initial phase of our experiment, we selected a random sample of 1000 sentences from the training set of the CONLL2003 dataset. This subset serves as the basis for assessing the annotation capabilities of LLM, aiming to explore the model’s proficiency in generating annotations comparable to the dataset’s meticulously crafted tags. Following the annotation process, the distilled data, enriched with LLM’s insights, along with another randomly selected sample of 1000 sentences from the original training set, are utilized in a mixed training approach to fine-tune a smaller model. The effectiveness of this strategy is subsequently evaluated on the CONLL2003 test set, allowing us to measure the impact of leveraging LLM annotations on model performance in named entity recognition tasks.
LLM Annotation
We employed GPT-4, specifically the gpt-4-0125-preview model provided by OpenAI, as our chosen Large Language Model (LLM) for annotating a subset of 1000 randomly sampled sentences from the CONLL2003 dataset. GPT-4 represents a significant leap forward in the domain of AI-driven natural language understanding and generation. As a large multimodal model, it has demonstrated unparalleled proficiency in accepting both text and image inputs, outputting text that showcases broader general knowledge and more advanced reasoning capabilities than its predecessors.
For our annotation process, we leveraged the GPT-4 turbo API to meticulously generate named entity annotations for the selected CONLL2003 data samples. This process involved employing two distinct prompting methods (detailed in subsequent sections), which facilitated the model’s understanding and execution of the annotation task. While GPT-4’s outputs were predominantly accurate and aligned with the task requirements, a minimal amount of manual post-processing was necessary to rectify occasional irregularities in the generated annotations, ensuring the highest quality of data for our study.
Standard Prompting
We employed a few-shot prompting approach, shown in Figure 1(left), with GPT-4 to annotate a random sample of 1000 sentences from the CONLL2003 dataset, serving as a baseline for comparison. Few-shot learning, a method where the model learns to perform tasks from a limited number of examples, was utilized to guide GPT-4 in accurately identifying and classifying named entities into four categories: LOC (Location), MISC (Miscellaneous), ORG (Organization), and PER (Person).
The prompt designed for this task succinctly introduced GPT-4 to its role as an NLP annotation expert, tasked with determining the type of each named entity present in a given sentence. It outlined the definitions for the four entity types, emphasizing the importance of showing the reasoning process in steps before presenting the results in a structured format. This approach aimed to leverage GPT-4’s advanced reasoning capabilities and vast general knowledge for generating precise entity annotations.
Included in the prompt were several example sentences, each followed by their expected annotation results, showcasing how entities should be categorized according to the predefined types. This structured prompt not only directed GPT-4 on how to approach the annotation task but also provided clear examples of the task’s expected outcomes, facilitating a more accurate and efficient annotation process.
CoT Prompting
Alongside the standard prompting approach, we implemented the Chain of Thought (CoT) prompting method combined with few-shot learning to annotate the same 1000 randomly sampled sentences from the CONLL2003 dataset using GPT-4, which is shown in Figure 1(right), . The CoT methodology is designed to mimic human-like reasoning by guiding the model through a stepwise thought process before arriving at a conclusion. This approach not only aims to improve the accuracy of the annotations by making the model’s thought process explicit but also seeks to provide insights into the model’s reasoning, thereby enhancing the interpretability of the annotations.
The CoT prompt we developed instructs GPT-4 to follow a four-step reasoning process for each sentence. It begins with understanding the context of the sentence, identifying potential named entities, determining the type of each entity based on context, and justifying the classification with reasoning. This structured approach encourages GPT-4 to articulate its thought process at each step, culminating in the structured format of named entity recognition results.
For instance, in annotating a sentence about a baseball game, GPT-4 is guided to recognize and classify entities such as "Houston" as a location (LOC) and "Orlando Miller" and "Todd Stottlemyre" as persons (PER), with justification provided for each classification. This not only aligns with the task’s requirements but also allows for an examination of how GPT-4 applies general knowledge and contextual reasoning to solve specific NLP tasks.
The use of CoT in conjunction with few-shot examples serves to further anchor GPT-4’s responses within the task’s framework, leveraging both the model’s advanced language understanding capabilities and its ability to engage in complex reasoning. By comparing the annotations generated through standard prompting with those produced via the CoT method, we aim to assess the impact of explicit reasoning on the quality and reliability of LLM-generated annotations for named entity recognition tasks.
Figure 1: The Standard Prompting Template (left) and CoT prompting Template (right)
Results and Analysis
Upon employing two distinct prompting strategies for GPT-4 annotation on a subset of 1000 sentences from the CONLL2003 dataset, our analysis yielded insightful results. The implementation of the Chain of Thought (CoT) method, which promotes an articulated step-by-step reasoning process, has demonstrated a pronounced advantage in performance over the standard few-shot approach. The results, as depicted in Table 1, indicate that the CoT prompting method achieved an overall F1-score of 0.73, which is superior to the 0.65 F1-score obtained using the standard few-shot prompting method. More specifically, the CoT method showed considerable improvements in precision and recall across most named entity types, with the most notable enhancement seen in the MISC category, where the CoT method’s precision more than doubled that of the few-shot method. This comparative performance analysis underscores the effectiveness of the CoT prompting method in mitigating the hallucination issue typically associated with LLMs. By prompting GPT-4 to engage in a transparent reasoning process, CoT prompting helped to align the model’s outputs more closely with the expected annotations, thereby improving the quality and reliability of the annotations for NER tasks. The higher performance metrics associated with CoT prompting suggest that guiding LLMs through a structured reasoning pathway can significantly amplify their reasoning abilities in complex NER tasks. This finding is pivotal, as it not only reinforces the value of CoT in enhancing the interpretability of LLM outputs but also demonstrates the potential for more accurate NER annotations, which is critical for downstream applications that rely on precise entity recognition.
Metric Standard Prompting CoT Prompting
Overall Precision 0.56 0.67
Recall 0.78 0.82
Micro Average F1 0.65 0.73
Macro Average F1 0.67 0.72
Weighted Average F1 0.73 0.77
LOC Precision 0.73 0.81
Recall 0.89 0.88
F1-Score 0.80 0.84
ORG Precision 0.73 0.77
Recall 0.67 0.76
F1-Score 0.70 0.77
PER Precision 0.97 0.95
Recall 0.96 0.96
F1-Score 0.97 0.95
MISC Precision 0.13 0.23
Recall 0.43 0.52
F1-Score 0.20 0.31
Table 1: Performance comparison between Standard and CoT Prompting
Phase Two: Training with Distilled Data
Model Introduction
In the subsequent phase of our research, we turned to the renowned BERT base model, specifically the uncased variant (bert-base-uncased), for fine-tuning with the data annotated by GPT-4. BERT, or Bidirectional Encoder Representations from Transformers, has been a transformative force in NLP, known for achieving state-of-the-art results across various tasks. Its architecture employs 12 layers of bidirectional transformers, making it adept at understanding the nuances of language context.
BERT’s uncased model, which was pre-trained on a large corpus, including the entirety of the English Wikipedia and the BookCorpus dataset, using a self-supervised learning approach. This approach involved two pre-training tasks: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), enabling the model to learn a deep bidirectional representation of the English language.
For our experiment, BERT’s bidirectional training framework makes it an ideal candidate to further distill and refine the knowledge transferred from LLMs like GPT-4. Its capacity to handle context effectively and its relatively compact architecture make it well-suited for scenarios with limited computational resources or closed-network environments, where leveraging the vast knowledge distilled from larger models like GPT-4 could be significantly beneficial.
The model we employed was structured to perform token classification, a fundamental task for NER, equipped with a classifier layer on top of the pre-trained BERT architecture. The use of BERT in this phase underscores our goal to capture the essence of the knowledge contained within the LLM-generated annotations and to utilize it in enhancing the model’s capability in a focused, resource-efficient manner.
Training with Original Data
We conducted a control experiment to establish a performance baseline for the BERT model. This was crucial for later comparison with the results obtained from the BERT model trained on the GPT-4 annotated data. For this purpose, we used another random subset of 1000 sentences from the CONLL2003 training set, excluding the portion annotated by GPT-4.
The intent behind using only the original dataset for training was to isolate the effect of the distilled data on the model’s learning capabilities. Our training parameters were meticulously set to optimize the learning process: the maximum sequence length was capped at 128 tokens, with a training batch size of 4, a validation batch size of 2, for a total of 10 epochs. The learning rate was set at 1e-05, with a max gradient norm of 10 to stabilize the training process.
Upon training, the model was evaluated on the CONLL2003 test set. The results showcased the inherent capability of the BERT model to capture and classify named entities, with an overall precision of 0.80, recall of 0.83, and an F1-score of 0.81. The evaluation scores across individual entity types—LOC, MISC, ORG, and PER—provided a granular view of the model’s performance, with the PER category achieving the highest F1-score of 0.92.
Training with Distilled Data Followed by Original Data
In refining our experiment’s second phase, we conducted a nuanced analysis of BERT’s performance when trained on a combination of distilled and original data. This phase was structured into two distinct training and evaluation steps to assess the efficacy of the distilled data from GPT-4 annotations in enhancing the model’s learning process.
Step 1: Initial Training with Distilled Data. BERT’s first training cycle utilized the 1000 sentences annotated by GPT-4. Despite employing consistent hyperparameters with the control experiment, the evaluation on the CONLL2003 test set resulted in a precision of 0.46, recall of 0.76, and an F1-score of 0.57. A closer look at the performance across different entity types revealed that the model achieved an F1-score of 0.76 for LOC, 0.14 for MISC, 0.65 for ORG, and 0.93 for PER, as depicted in the provided evaluation details. These scores indicated a substantial variance across categories, with particularly lower scores in the MISC category, suggesting that the distilled data alone may not provide enough variety for the model to generalize effectively across all entity types.
Step 2: Augmented Training with Original Data. Subsequent to the initial training, BERT was further trained on an additional 1000 sentences randomly sampled from the CONLL2003 training dataset. This dataset was identical to the one used in the control group to maintain consistency and ensure a meaningful comparison. After completing the second training cycle with the same hyperparameters and an additional 10 epochs, BERT’s performance on the test set improved significantly, achieving an overall precision of 0.83, recall of 0.87, and an F1-score of 0.85. The detailed scores for individual entity types were 0.88 for LOC, 0.60 for MISC, 0.84 for ORG, and 0.95 for PER, reflecting considerable improvements across the board and notably addressing the earlier shortcomings in the MISC category.
Results and Analysis
The experimental results provided a compelling narrative on the effectiveness of using a combined dataset for training the BERT model (shown in Table 2). When the model was initially trained solely on the distilled dataset annotated by GPT-4, the overall test performance yielded an F1-score of 0.57, indicating that while the model could capture the essence of the data for certain entity types, the overall effectiveness was not optimal. Notably, the performance on individual entity types such as LOC and PER was relatively strong, with F1-scores of 0.76 and 0.93 respectively, suggesting that GPT-4’s annotations were particularly beneficial for these categories.
However, the MISC category presented a significant challenge, with an F1-score of only 0.14. This could be attributed to the model’s unfamiliarity with the diverse and less defined nature of the MISC entities, compounded by GPT-4’s inclination to generalize unfamiliar content under the MISC label, which adversely affected the overall metrics.
In contrast, when BERT was subsequently trained on a combination of distilled and original data, the model’s performance improved dramatically, achieving an F1-score of 0.85 in the final test evaluation. This improvement was consistent across all entity types, with even the MISC category showing a notable increase to an F1-score of 0.60. This two-step training process, using both GPT-4 annotated data and human-annotated data, proved to be more effective than using either dataset alone.
The results strongly suggest that leveraging distilled data from advanced LLMs like GPT-4, when complemented with original human annotations, can significantly enhance the performance of NLP models. This strategy effectively combines the broad contextual knowledge of LLMs with the nuanced understanding of human annotators, leading to a more robust and efficient named entity recognition capability. Such an approach offers a promising avenue for reducing costs and increasing efficiency in NLP applications, validating the integration of LLM insights into traditional NLP tasks.
Metric Pure CONLL Data Training Distilled Data Training Combined Data Training
Overall
   Precision 0.80 0.46 0.83
   Recall 0.83 0.76 0.87
   F1-Score 0.81 0.57 0.85
LOC
   Precision 0.78 0.70 0.90
   Recall 0.89 0.84 0.86
   F1-Score 0.83 0.76 0.88
MISC
   Precision 0.52 0.08 0.52
   Recall 0.67 0.54 0.70
   F1-Score 0.59 0.14 0.60
ORG
   Precision 0.80 0.69 0.81
   Recall 0.77 0.62 0.86
   F1-Score 0.78 0.65 0.84
PER
   Precision 0.96 0.95 0.94
   Recall 0.89 0.91 0.96
   F1-Score 0.92 0.93 0.95
Table 2: Comparison of BERT model performance trained on pure CONLL data, distilled data, and a combination of both
Discussion
In the discussion section of our paper, we delve into the merits and limitations of our proposed solution, as well as outline avenues for future work.
Advantages of the Proposed Solution
Our approach reduces the need for extensive human annotation, thereby increasing efficiency. The results demonstrate that models trained on a mix of GPT-4 annotated data and human annotations outperform those trained on human annotations alone. This method is particularly advantageous for cold-start scenarios, where using GPT-4 for pre-annotation or training models on pre-annotated data provides them with a foundational capability, aiding annotators in quickly ramping up the annotation process.
The hybrid training of small models with a mix of distilled data and human annotations presents higher adaptability for scenarios with limited resources or closed network environments. This adaptability is critical for applications where large language models cannot be directly accessed.
The Chain of Thought (CoT) methodology enhances the interpretability of the annotation process, allowing for easier validation and corrections of the results. The detailed reasoning provided by CoT prompts not only improves the inference performance but also serves an educational purpose, guiding novice annotators in the tagging process.
The use of structured prompts outputting dictionaries, such as ’LOC’: [’Houston’], ’PER’: [’Orlando Miller’, ’Todd Stottlemyre’], ’ORG’: [’Houston Astros’, ’St. Louis Cardinals’, ’NL Central division’], enables the efficient representation of all entity types in a single request. This format can easily handle non-continuous entities and nested entities (e.g., "national basketball of America" as an made up organization containing "America" as a location), and it facilitates automated processing through conversion to dictionary objects. Additionally, the reasoning process uncovers certain annotation issues within the CONLL dataset, where GPT’s classifications sometimes appear more logical than the original labels.
Limitations
The output from LLMs like GPT-4 can still be unstable, occasionally deviating from the desired format or presenting entities that do not match the source text exactly, such as inconsistencies in capitalization or content "corrections" based on the model’s interpretation. This necessitates manual review and correction.
The hallucination issue in GPT models persists, with a tendency to over-identify entities, which can reduce the accuracy of the MISC category.
One potential limitation is the dependency on the quality and the relevance of the few-shot examples provided to GPT-4. If the examples are not closely aligned with the target domain or context, the quality of the distilled data may suffer.
Future Work
To address the first limitation, future research could explore more advanced LLMs or optimize the phrasing of prompts. Additionally, parameterized prompting could be considered for LLMs deployable in a localized setting.
For the second limitation, we could refine our approach by selecting few-shot examples based on embedding similarity, dynamically choosing examples that are more closely related to the entity types in the target annotation task, thereby improving accuracy.
Implementing a tiered selection process for samples annotated by LLMs might be beneficial. High-quality LLM annotations could be identified and used for further pre-training of smaller models, enhancing their performance.
Integrating active learning strategies where LLMs focus on annotating more challenging samples could boost the efficiency of the smaller model’s learning process.
Ongoing optimization of training strategies, exploring new combinations of LLMs and smaller models, and extending the approach to a wider array of NLP tasks are all promising directions for future research.
Conclusion
In conclusion, our research corroborates the significant potential of integrating Large Language Models (LLMs) such as GPT-4 into traditional NLP tasks, particularly Named Entity Recognition (NER). By employing the Chain of Thought (CoT) technique, we have successfully distilled knowledge from LLMs and utilized it to enhance the training of smaller models like BERT.
Our mixed-training strategy, which combines LLM-generated annotations with human annotations, has proven to be more effective than reliance on human efforts alone. This approach not only improves efficiency and performance but also reduces costs, aligning with our initial aim to leverage the strengths of LLMs while addressing their limitations in capacity and hallucination.
The application of our method is especially pertinent in resource-constrained and closed-network environments, showcasing the adaptability and practicality of smaller models when large-scale LLMs are not feasible.
Reflecting on our discussion, we recognize the challenges of LLM variability and the ongoing issues with hallucination. Future work will focus on advancing the sophistication of prompts, exploring parameterized prompting, and refining selection processes for LLM annotations to further improve the quality of NER models. Our continued efforts aim to optimize training strategies and expand the reach of this methodology across diverse NLP tasks.